OpenStudy (anonymous):

Let's let the equation y = 0.50(1.06)x represent the price of movie tickets for the years after 1950. You have filled in the ticket table for the prices for the years 0, 10, 20, 30, 40. Do these prices look reasonable?

OpenStudy (anonymous):

# years after 1950 X 0 10 20 30 40 Ticket Price $ Y " Facts: y = 0.50(1.06)x X = 0, 10, 20, 30, and 40 after 1950, 1960, 1970, 1980, and 1990 y= .50 times 1.06 times x for each year after 1950 x represents the number you are going to multiply regarding the year, so for example for 1960 the equation would look like this y = .50*1.06*10 1970 y = .50*1.06*20 1980 y = .50*1.06*30 and so on..... Do these prices look reasonable? yes